(3x-1)^2=4/9. Solve using square root property
Answer by stanbon(75887)

(3x-1)^2=4/9. Solve using square root property
----------------
3x-1 = 2/3 or 3x-1 = -2/3
3x = 5/3 or 3x = 1/3
---
x = 5/9 or x = 1/9
=====================
Cheers,
Stan H.